Consider an RLC series circuit with R = 600 Ω, L = 3 H, C = 4μF, generator voltage V = 20 v, frequency= 60 hz.

Find

a) the inductive impedance XL,

b) capacitive impedance Xc ,

c) Total impedance Z,

d) Line current I ,

e) Voltage drops VR , VL, ,Vc

f) combination voltage VRL , and VLc ,

g) phase angle φ ,

h) resonant frequency f0 ,

i) Power dissipated by circuit.
Answer #1

Giveno- R=gooch L = 3H C = 4uF = 4x166F V - 20 Volt f = 60 Hz We know that, Angular frequency (W) = 217f - 2x Tx60 = 376.993(d) Line Current (i) - ė = V z = 20 760.06 i = 0.0263 A ANS (e) Voltage Drops- è R V R = 0.0263x600 VR = 15.78 V. ANS V i XLVL P(VL-Vc) And Kc = - Ve = 29.75-17.44 Vc Vis = 12-31 Volt/ ANS Phase diagram 19) Phase angle (0) VL-Ve VR 29.75-17.44 15.7lis power dissipated by circuit (P) – P=vx = 20 x 0.0263 Р =0.526 Walt | ANS